A satellite of mass 200 kg is displaced by a distance of



2RE



from a point on the



Earth’s surface, where



RE



is the radius of the Earth. Calculate the work done on



the satellite by the force of gravitation. Take the values of the constants in your



calculation from the astrophysical data given in the course

1
Expert's answer
2022-02-28T10:05:48-0500

"h=2R,"

"U=\frac 12 mg'h,"

"g'=g\frac{R^2}{(R+h)^2},"

"U=\frac{mgR^2\cdot 2R}{2(R+2R)^2}=\frac{mgR}{9}=1.4~GJ."